Circular No. 6030

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) D. C. Jewitt and D. J. Tholen, University of Hawaii, succeeded in

securing positions of nuclei P_{2} = 8b, Q_{1} = 7a, Q_{2} = 7b, R = 6, S = 5 and W = 1 with the 2.2-m reflector at Mauna Kea on July 19.3 UT, and new orbit determinations by P. W. Chodas and D. K. Yeomans, Jet Propulsion Laboratory, have refined the impact times on Jupiter as follows: P_{2}, July 20.636 UT; Q_{2}, 20.824; Q_{1}, 20.836; R, 21.228; S, 21.634; W = 22.333. Observations of the impacts that have already occurred indicate that the predictions on IAUC 6022 tended to be up to about 0.007 day too early, and it is possible that the new predictions will still be too early, a problem that can be attributed to systematic errors in the Guide Star Catalogue (cf. IAUC 6017).

Among the wealth of messages received via the SL9 message center, one from J. Klavetter indicates that the impact of the missing fragment M = 10 (cf. IAUC 6017) was observed at the San Pedro Martir station of the Mexican National Astronomical Observatory. Observations mainly at 2.1 μm showed a dim spot just beyond site C on July 20.259 UT, and it continued to brighten until 20.299.

D. Crisp reports for the IRIS observers on the 3.9-m Anglo-Australian Telescope that the impact flash from fragment N = 9 was detected, using rapid-sampling imaging photometry with a 2.34-μm filter rather than Kgrism drift scanning, on July 20.441 UT. The intensity of this flash increased slowly until about 20.442, but it never exceeded 70 percent of the mean surface brightness of the south polar hood. On resuming the observations on 20.446 after a problem with the instrument there was no evidence of the N impact site. K-grism drift scanning from July 20.451 to 20.458 also showed no evidence of the N impact site. P. McGregor and M. Allen, Australian National Observatory, report that CASPIR 2.34-μm observations showed the N impact as a faint flash beginning on July 20.441 and lasting until 20.443, but subsequent narrow-band imaging failed to detect the site.

Following up their earlier report (IAUC 6028), Lellouch et al. have continued to monitor CO emission at 230-GHz with the IRAM 30-m telescope. The detection at impact sites G and H was confirmed on July 19.8 UT at integrated intensities 1.6 and 1 K km/s, respectively. Re-examination of data from July 17.8 UT showed the emission at 0.6 K km/s at site E and marginally at site C, and a marginal detection from site E was also made on July 19.8. The width (FWHM) of the clearly detected lines is constant at 2–3 km/s. Unsuccessful searches were made for CO^{+} and H_{2}S.

1994 July 20 Brian G. Marsden